# Speaking at the Right Level: Literacy-Controlled Counterspeech Generation with RAG-RL

### v2

@@ 1 Introduction
Social media users actively participate in combating health misinformation by posting counterspeech Kuo and Chen (2025); Anik et al. (2025); Micallef et al. (2020). However, user-generated responses often lack credible evidence, decreasing the effectiveness of counterspeech Yue et al. (2024a). While experts and fact-checkers can provide factual support in debunking health misinformation, it becomes less efficient when confronted with an increasing volume of health misinformation Yue et al. (2024b). Thus, generating supportive counterspeech at scale remains a challenge.
Recent studies have explored the use of Large Language Models (LLMs) to generate counterspeech in response to health misinformation Yue et al. (2024a, b); Wang et al. (2024b); He et al. (2023). However, these methods often overlook a critical issue: Is the generated counterspeech accessible and persuasive to users with different health literacy levels? Health literacy refers to the capability to understand, judge, and use health information to make good choices and stay healthy World Health Organization (2024). The health literacy level of the audience can significantly affect the effectiveness of counterspeech Liu et al. (2020). Users with lower health literacy often struggle to comprehend counterspeech grounded in scientific research Shahid et al. (2022); Liu et al. (2020) and understand information that contains complex or technical language Centers for Disease Control and Prevention (2024); Chen et al. (2018). In contrast, users with a higher literacy level may lose interest or perceive the content as oversimplified if it is tailored to a low comprehension level August et al. (2024); Mart´ınez Silvagnoli et al. (2022). For example, simplistic explanations such as “It is made from a tiny, weak part of the virus that cannot make you sick …” in Figure 1 (low-level counterspeech) may be inappropriate for users with advanced health literacy. Such mismatches in communication can critically undermine the effectiveness of counterspeech.
@@ 2.1 Counterspeech to Misinformation Generation
Counterspeech has been proven effective in mitigating misinformation Peng and Grimmelmann (2024); Siebert and Siebert (2023). Previous studies focus on generating counterspeech with desirable attributes, such as evidence-based content, factual accuracy, persuasiveness, or overall effectiveness Anik et al. (2025); Yue et al. (2024a); Hong et al. (2024); Yue et al. (2024b); He et al. (2023). For example,  Anik et al. (2025) and Yue et al. (2024a) combined external scientific sources to generate evidence-based counterspeech in refuting misinformation.  Yue et al. (2024b) generated factually accurate counterspeech by synthesizing contrastive arguments derived from fact-checking sources.  He et al. (2023) proposed a reinforcement learning-based framework to generate counterspeech with enhanced politeness, factuality, and refutational strength. Despite these efforts, existing studies have not yet investigated counterspeech generation tailored to users with different health literacy levels.
@@ 4.3 Experiment Setup
We experiment with various LLMs, including models with similar parameter sizes but different architectures (e.g., LLaMA3.1-8B-Instruct [^2] Available at: https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct vs. Qwen2.5-7B-Instruct [^3] Available at: https://huggingface.co/Qwen/Qwen2.5-7B-Instruct), and models from the same family with varying parameter sizes (e.g., LLaMA3.1-8B-Instruct vs. LLaMA3.2-1B-Instruct [^4] Available at: https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct)). This setup allows us to investigate model performance across diverse architectures and parameter scales.
@@
RL Optimization We fine-tune LLMs using GRPO to generate counterspeech tailored to users with low, medium, and high health literacy. The policy backbone is a LoRA-adapted supervised model, initially fine-tuned on the MisinfoCorrect dataset, which pairs health misinformation with corresponding counterspeech to prevent response refusal when prompted with misinformation. For RL, we use the MisinfoLiteracy dataset, splitting it into 80% for GRPO training and 20% for inference evaluation. The reward function combines three components: (1) a double-sigmoid FKRE-based score that promotes readability within a specific target range (e.g., 80–100 for low literacy), and (2) a GPT-4o-mini-based user preference score, rated on a 1–5 Likert scale, measuring how well the generated counterspeech aligns with user expectations. (3) These two signals are weighted equally (0.5 readability, 0.5 user preference) and aggregated to compute the total reward. For each prompt, LLMs generate four responses and compute their relative advantages based on an aggregated reward signal. The GRPO training loop then updates the policy to increase the likelihood of higher-reward responses while constraining divergence from a fixed reference policy (initialized from the supervised LoRA-adapted model) This setup enables the model to learn responses that are both accessible and effective across varying levels of health literacy. The GRPO training is configured with the following hyperparameters: a per-device batch size of 4, gradient accumulation steps of 1, a learning rate of $5\times 10^{-6}$, and training epochs of 3. For each prompt, the model generates 4 completions with a maximum token length of 200. The GRPO-specific regularization coefficient $\beta$ is set to 0.2.
@@ 4.4 Evaluation
User Preference This dimension captures how a response is subjectively perceived by users with different health literacy levels. It offers additional insight into how well a counterspeech message connects with its intended audience, for example, how people with low literacy levels view the effectiveness of messages designed for them, as well as those designed for medium or high literacy audiences. Due to the high cost of recruiting real users with diverse literacy levels, we employ LLMs as evaluators. These LLMs are guided with tailored instructions to simulate users at specific health literacy levels. We use a 1–5 Likert-style scale to assess user preference, with detailed evaluation prompts documented in the Appendix D.2. We further conduct human validation for the user preference evaluation (see details in Appendix E). Before the evaluation, we administer a brief health literacy survey adapted from Rasmussen et al. (2023), consisting of several short questions to assess participants’ health literacy levels. Based on the survey results, we recruit participants representing low, medium, and high health literacy groups to ensure diversity and alignment with target user profiles. The results in Table LABEL:tab:human_validation show that the agreement rate among human annotators exceeds 0.85, with Cohen’s Kappa values above 0.60, indicating substantial agreement. The agreement rate between human judgments and LLM-generated evaluations is above 0.90, and the Cohen’s Kappa exceeds 0.60, indicating a reliable evaluation by the LLM evaluator. In addition, we have conducted a correlation analysis between human ratings and LLM-generated ratings of user preference (see Table LABEL:tab:correlation ). The results suggest that the LLM’s behavior adapts well to different user profiles. We document the analysis in Appendix E Correlation Analysis.
@@ 4.5 Results
We present the evaluation of results in Table LABEL:tab:Main_results. There are several findings we conclude as follows.
@@ 5 Cross Evaluation
Controlled-Literacy generalizes robustly across datasets. We report the results of the first evaluation in Table LABEL:tab:checkcovid_results and Table LABEL:tab:misinfo_correct_results in Appendix G. Across the CheckCovid and MisInfoCorrect datasets, the Controlled-Literacy method consistently achieves the best average performance in all four metrics: (1) Politeness: Highest across nearly all model configurations (e.g., 0.85 in LLaMA-8B for CheckCovid, 0.86 in LLaMA-8B for MisInfoCorrect). (2) Target Distance: Consistently the lowest among all baselines, indicating superior alignment with the intended readability level. (3) User Preference: Typically equal to or higher than other methods (e.g., 0.75 in most configurations). (4) Factual accuracy: Always higher than RAG and Instructional Prompt, which indicates that counterspeech generated by Controlled-Literacy is more reliable.
@@
User preference peaks when counterspeech matches literacy level. The second evaluation results in Table LABEL:tab:Cross_evaluation show that the top preference scores for each user group generally occur when the counterspeech matches the user’s literacy level. For instance, in the Instructional Prompt setting, low and high users show the highest preference for low-level (0.75) and high-level (0.75) responses, respectively. Similarly, under the RAG setting, the top scores appear for low (0.73), medium (0.73), and high (0.75) users when aligned with their corresponding levels. The Controlled-Literacy method shows a similar trend, with low users preferring low-level counterspeech (0.74) and high users favoring high-level responses (0.75).
However, there are notable exceptions where users prefer counterspeech designed for different literacy levels. For example, in the Instructional Prompt setting, medium users rated low-level counterspeech slightly higher (0.75) than medium-level (0.74). In the RAG setting, medium users rated both low and medium-level counterspeech equally (0.73). Likewise, in the Controlled-Literacy setting, medium users showed a higher preference for low-level responses (0.74) compared to medium-level (0.73). These findings suggest that while aligning counterspeech with user literacy level generally yields the best user preference outcomes, users may sometimes prefer responses with a lower readability level, possibly due to greater clarity or accessibility. This insight inspires us to further optimize our generation strategy by aligning responses more closely with the lower end of the target health literacy range in the future.
@@ 6 Qualitative Analysis
They summarize several key elements that make the counterspeech achieve high performance: (1) Evidence-Based reasoning: most of the counterspeech cite scientific sources (CDC, WHO, peer-reviewed studies) and explain the details, which matches the expectations of rational justification for high health literacy level of users. (2) Precise terminology. The counterspeech uses domain-specific vocabulary, such as “pathogenesis,” “clinical trials" and so on. (3) Structured argumentation. The counterspeech often follows a consistent structure: Acknowledgement → Clarification → Evidence → Implication → Recommendation, satisfying expectations for logical coherence and depth.
However, our best-performing responses also exhibit several shortcomings that need further improvement: (1) Repetitive openings. Many counterspeech begin with similar phrases (e.g., "Thank you for sharing…"), which lack lexical variety. It tends to be less human-like, potentially decreasing the effectiveness of counterspeech. (2) Heavy information density. While users with high health literacy generally prefer more technical content, some responses are overly dense and packed with information. It may diminish clarity and fatigue readers. (3) Detached tone. Although the counterspeech maintains a coherent and logically structured academic tone, it often lacks emotional resonance or personal storytelling. Incorporating more empathetic or narrative elements could enhance reader engagement and strengthen the persuasive impact.

### v3

@@ 1 Introduction
Social media users actively participate in combating health misinformation by posting counterspeech Kuo and Chen (2025); Anik et al. (2025); Micallef et al. (2020). However, user-generated responses often lack credible evidence, decreasing the effectiveness of counterspeech Yue et al. (2024a). While experts and fact-checkers can provide factual support in debunking health misinformation, it becomes less efficient when confronted with an increasing volume of health misinformation Yue et al. (2024b). Generating supportive counterspeech at scale remains a challenge.
Recent studies have explored the use of Large Language Models (LLMs) to generate counterspeech in response to health misinformation Yue et al. (2024a, b); Wang et al. (2024b); He et al. (2023). However, these methods often overlook a critical issue: Is the generated counterspeech accessible and persuasive to users with different health literacy levels? Health literacy refers to the capability to understand, judge, and use health information to make good choices and stay healthy World Health Organization (2024). The health literacy level of the audience can significantly affect the effectiveness of counterspeech Liu et al. (2020). Users with lower health literacy often struggle to comprehend counterspeech grounded in scientific research Shahid et al. (2022); Liu et al. (2020) and understand information that contains complex or technical language Centers for Disease Control and Prevention (2024); Chen et al. (2018). In contrast, users with a higher literacy level may lose interest or perceive the content as oversimplified if it is tailored to a low comprehension level August et al. (2024); Martínez Silvagnoli et al. (2022). For example, simplistic explanations such as “It is made from a tiny, weak part of the virus that cannot make you sick …” in Figure 1 (low-level counterspeech) may be inappropriate for users with advanced health literacy. Such mismatches in communication can critically undermine the effectiveness of counterspeech.
@@ 2.1 Counterspeech to Misinformation Generation
Counterspeech has been proven effective in mitigating misinformation Peng and Grimmelmann (2024); Siebert and Siebert (2023). Previous studies focus on generating counterspeech with desirable attributes. For example, Anik et al. (2025) and Yue et al. (2024a) combined external scientific sources to generate evidence-based counterspeech in response to misinformation.  Yue et al. (2024b) generated factually accurate counterspeech by synthesizing contrastive arguments derived from fact-checking sources.  Hong et al. (2024) adapted the generation of counterspeech according to preferred conversation outcomes, resulting in more positive conversation engagement.  He et al. (2023) proposed a reinforcement learning-based framework to generate counterspeech with enhanced politeness, factuality, and refutational strength. Despite these efforts, existing studies have not yet investigated counterspeech generation tailored to users with different health literacy levels.
@@ 4.2 Baseline
Method Literacy Level Politeness Target Distance (↓) User Preference Factual Accuracy
LLaMA-8B
Instructional Prompt low 0.39 (0.23) 2.16 (3.69) 0.75 (0.03) 0.86
medium 0.49 (0.22) 5.99 (8.24) 0.74 (0.06) 0.89
high 0.35 (0.19) 0.07 (0.74) 0.75 (0.04) 0.87
Avg. 0.41 (0.21) 2.74 (4.22) 0.75 (0.04) 0.87
RAG low 0.72 (0.25) 1.30 (3.23) 0.71 (0.13) 0.89
medium 0.66 (0.22) 4.45 (6.35) 0.71 (0.13) 0.88
high 0.41 (0.21) 0.08 (0.75) 0.72 (0.09) 0.90
Avg. 0.60 (0.23) 1.94 (3.44) 0.71 (0.12) 0.89
Controlled-Literacy low 0.84 (0.15) 1.21 (2.21) 0.74 (0.07) 0.89
medium 0.69 (0.20) 1.50 (3.81) 0.73 (0.10) 0.90
high 0.98 (0.02) 0.00 (0.00) 0.75 (0.00) 0.93
Avg. 0.84 (0.15) 0.90 (2.01) 0.74 (0.04) 0.91
Qwen-7B
Instructional Prompt low 0.36 (0.22) 0.37 (3.54) 0.74 (0.06) 0.71
medium 0.48 (0.23) 4.25 (6.50) 0.73 (0.13) 0.88
high 0.41 (0.18) 0.37 (3.54) 0.73 (0.12) 0.89
Avg. 0.42 (0.15) 1.66 (4.53) 0.73 (0.09) 0.83
RAG low 0.69 (0.25) 0.60 (1.97) 0.69 (0.16) 0.79
medium 0.52 (0.18) 3.07 (5.55) 0.68 (0.18) 0.92
high 0.45 (0.19) 0.01 (0.13) 0.73 (0.07) 0.91
Avg. 0.55 (0.13) 1.23 (2.55) 0.70 (0.13) 0.87
Controlled-Literacy low 0.77 (0.18) 2.09 (3.38) 0.73 (0.11) 0.84
medium 0.55 (0.19) 4.08 (5.67) 0.74 (0.08) 0.94
high 0.86 (0.22) 0.00 (0.02) 0.75 (0.04) 0.92
Avg. 0.73 (0.16) 2.06 (3.02) 0.74 (0.08) 0.90
LLaMA-1B
Instructional Prompt low 0.65 (0.28) 8.17 (9.36) 0.66 (0.17) 0.50
medium 0.40 (0.25) 17.30 (23.39) 0.58 (0.22) 0.58
high 0.70 (0.24) 0.19 (2.87) 0.68 (0.14) 0.54
Avg. 0.58 (0.16) 8.55 (11.87) 0.64 (0.17) 0.54
RAG low 0.50 (0.31) 16.90 (14.34) 0.45 (0.25) 0.63
medium 0.41 (0.21) 9.64 (12.20) 0.36 (0.25) 0.71
high 0.42 (0.22) 0.51 (4.92) 0.42 (0.22) 0.63
Avg. 0.44 (0.05) 9.02 (10.49) 0.34 (0.24) 0.66
Controlled-Literacy low 0.73 (0.17) 2.08 (3.62) 0.68 (0.15) 0.61
medium 0.63 (0.21) 1.86 (4.10) 0.71 (0.12) 0.75
high 0.85 (0.27) 0.00 (0.00) 0.67 (0.13) 0.76
Avg. 0.74 (0.11) 1.31 (2.57) 0.69 (0.13) 0.71
Table 1: Counterspeech generation results on MisinfoLiteracy categorized by health literacy levels: low, medium, and high, along with the overall average. The mean(variance) across samples is reported for Politeness, Target Distance, and User Preference. Factual Accuracy reports the percentage of responses that are factually correct. User preference presents the evaluation by intended user category. Higher mean values for politeness, user preference, and factual accuracy indicate better performance, while lower mean values for target distance indicate better alignment. The best overall performance in each category is highlighted in gray.
@@ 4.3 Experiment Setup
We experiment with various LLMs, including models with similar parameter sizes but different architectures (e.g., LLaMA3.1-8B-Instruct [^2] Available at: https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct vs. Qwen2.5-7B-Instruct [^3] Available at: https://huggingface.co/Qwen/Qwen2.5-7B-Instruct), and models from the same family with varying parameter sizes (e.g., LLaMA3.1-8B-Instruct vs. LLaMA3.2-1B-Instruct [^4] Available at: https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct). This setup allows us to investigate model performance across diverse architectures and parameter scales.
@@
RL Optimization We fine-tune LLMs using GRPO to generate counterspeech tailored to users with low, medium, and high health literacy. The policy backbone is a LoRA-adapted supervised model, initially fine-tuned on the MisinfoCorrect dataset, which pairs health misinformation with corresponding counterspeech to prevent response refusal when prompted with misinformation. For RL, we use the MisinfoLiteracy dataset, splitting it into 80% for GRPO training and 20% for inference evaluation. The reward function combines two components: (1) a double-sigmoid FKRE-based score that promotes readability within a specific target range (e.g., 80–100 for low literacy), and (2) a GPT-4o-mini-based user preference score, rated on a 1–5 Likert scale, measuring how well the generated counterspeech aligns with user expectations. These two signals are weighted equally (0.5 readability, 0.5 user preference) and aggregated to compute the total reward. For each prompt, LLMs generate four responses and compute their relative advantages based on an aggregated reward signal. The GRPO training loop then updates the policy to increase the likelihood of higher-reward responses while constraining divergence from a fixed reference policy (initialized from the supervised LoRA-adapted model). This setup enables the model to learn responses that are both accessible and effective across varying levels of health literacy. The GRPO training is configured with the following hyperparameters: a per-device batch size of 4, gradient accumulation steps of 1, a learning rate of $5\times 10^{-6}$, and training epochs of 3. For each prompt, the model generates 4 completions with a maximum token length of 200. The GRPO-specific regularization coefficient $\beta$ is set to 0.2.
@@ 4.4 Evaluation
Literacy Level Category Tolerant Match Cohen’s Kappa
Low Computing 0.94 0.65
Human 0.96 0.75
Medium Computing 0.96 0.65
Human 0.92 0.69
High Computing 0.96 0.65
Human 0.88 0.63
Table 2: Agreement between human-based and computing-based evaluations and inter-annotator agreement among human evaluators on User Preference.
@@
Literacy Level Pearson Spearman Kendall’s Tau
Low level 0.68 0.65 0.59
Medium level 0.67 0.60 0.56
High level 0.61 0.73 0.70
Table 3: Correlation between human ratings and LLM-generated ratings across literacy levels
User Preference This dimension captures how a response is subjectively perceived by users with different health literacy levels. It offers additional insight into how well a counterspeech message connects with its intended audience, for example, how people with low literacy levels view the effectiveness of messages designed for them, as well as those designed for medium or high literacy audiences. Due to the high cost of recruiting real users with diverse literacy levels, we employ LLMs as evaluators. These LLMs are guided with tailored instructions to simulate users at specific health literacy levels. We use a 1–5 Likert-style scale to assess user preference, with detailed evaluation prompts documented in the Appendix D.2. We further conduct human validation for the user preference evaluation (see details in Appendix E). Before the evaluation, we administer a brief health literacy survey adapted from Rasmussen et al. (2023), consisting of several short questions to assess participants’ health literacy levels. Based on the survey results, we recruit participants representing low, medium, and high health literacy groups to ensure diversity and alignment with target user profiles. The results in Table 2 show that the agreement rate among human annotators exceeds 0.85, with Cohen’s Kappa values above 0.60, indicating substantial agreement. The agreement rate between human judgments and LLM-generated evaluations is above 0.90, and the Cohen’s Kappa exceeds 0.60, indicating a reliable evaluation by the LLM evaluator. In addition, we have conducted a correlation analysis between human ratings and LLM-generated ratings of user preference (see Table 3 ). The results suggest that the LLM’s behavior adapts well to different user profiles. We document the analysis in Appendix E Correlation Analysis.
@@ 4.5 Results
We present the evaluation of results in Table 1. There are several findings we conclude as follows.
@@ 5 Cross Evaluation
Counterspeech/User Low User Medium User High User
Instructional Prompt
   low 0.75 (0.03) 0.75 (0.04) 0.68 (0.12)
   medium 0.73 (0.08) 0.74 (0.06) 0.74 (0.07)
   high 0.55 (0.16) 0.62 (0.15) 0.75 (0.04)
RAG
   low 0.73 (0.09) 0.73 (0.10) 0.50 (0.17)
   medium 0.72 (0.11) 0.73 (0.11) 0.66 (0.17)
   high 0.45 (0.18) 0.51 (0.18) 0.72 (0.09)
Controlled-Literacy
   low 0.74 (0.07) 0.74 (0.05) 0.61 (0.16)
   medium 0.69 (0.16) 0.73 (0.10) 0.62 (0.23)
   high 0.62 (0.15) 0.63 (0.13) 0.75 (0.00)
Table 4: Cross evaluation of user preference.
Controlled-Literacy generalizes robustly across datasets. We report the results of the first evaluation in Table 8 and Table 9 in Appendix G. Across the CheckCovid and MisInfoCorrect datasets, the Controlled-Literacy method consistently achieves the best average performance in all four metrics: (1) Politeness: Highest across nearly all model configurations (e.g., 0.85 in LLaMA-8B for CheckCovid, 0.86 in LLaMA-8B for MisInfoCorrect). (2) Target Distance: Consistently the lowest among all baselines, indicating superior alignment with the intended readability level. (3) User Preference: Typically equal to or higher than other methods (e.g., 0.75 in most configurations). (4) Factual accuracy: Always higher than RAG and Instructional Prompt, which indicates that counterspeech generated by Controlled-Literacy is more reliable.
@@
User preference peaks when counterspeech matches literacy level. The second evaluation results in Table 4 show that the top preference scores for each user group generally occur when the counterspeech matches the user’s literacy level. For instance, in the Instructional Prompt setting, low and high users show the highest preference for low-level (0.75) and high-level (0.75) responses, respectively. Similarly, under the RAG setting, the top scores appear for low (0.73), medium (0.73), and high (0.75) users when aligned with their corresponding levels. The Controlled-Literacy method shows a similar trend, with low users preferring low-level counterspeech (0.74) and high users favoring high-level responses (0.75).
However, there are notable exceptions where users prefer counterspeech designed for different literacy levels. For example, in the Instructional Prompt setting, medium users rated low-level counterspeech slightly higher (0.75) than medium-level (0.74). In the RAG setting, medium users rated both low and medium-level counterspeech equally (0.73). Likewise, in the Controlled-Literacy setting, medium users showed a higher preference for low-level responses (0.74) compared to medium-level (0.73). These findings suggest that while aligning counterspeech with user literacy level generally yields the best user preference outcomes, users may sometimes prefer responses with a lower readability level, possibly due to better accessibility. This inspires us to further optimize our generation strategy by aligning responses more closely with the lower end of the target health literacy range in the future.
@@ 6 Qualitative Analysis
They summarize several key elements that make the counterspeech achieve high performance: (1) Evidence-based reasoning: most of the counterspeech cite scientific sources (CDC, WHO, peer-reviewed studies) and explain the details, which matches the expectations of rational justification for high health literacy level of users. (2) Precise terminology. The counterspeech uses domain-specific vocabulary, such as “pathogenesis,” “clinical trials” and so on. (3) Structured argumentation. The counterspeech often follows a consistent structure: acknowledgement → clarification → evidence → implication → recommendation, satisfying expectations for logical coherence and depth.
However, our best-performing responses also exhibit several shortcomings that need further improvement. (1) Repetitive openings. Many counterspeech begin with similar phrases (e.g., “Thank you for sharing…”), which lack lexical variety. It tends to be less human-like, potentially decreasing the effectiveness of counterspeech. (2) Heavy information density. While users with high health literacy generally prefer more technical content, some responses are overly dense and packed with information. It may diminish clarity and fatigue readers. (3) Detached tone. Although the counterspeech maintains a coherent and logically structured academic tone, it often lacks emotional resonance or personal storytelling. Incorporating more empathetic or narrative elements could enhance reader engagement and strengthen the persuasive impact.
@@ B.2 Human validation
Document Name Source
COVID-19 Activity Book Johns Hopkins Children’s Center Child Life Program
Covid-19-Plain-Language-Flyer-with-Facemask Independent Living Center of the Hudson Valley
COVID-19-Vaccines-A-Plain-Language-Guide HealthMatters Program
FINAL Diagnostic Testing U.S. Centers for Disease Control and Prevention
Healthy School Year U.S. Centers for Disease Control and Prevention
Pediatric Testing Materials Johns Hopkins Children’s Center Child Life Program
Plain Language COVID Fact Sheets Maryland Developmental Disabilities Council
Plain Language Guide for COVID-19_Group-Home St. Clair County Community Mental Health (SCCCMH)
COVID-19 Teen Info Sheet U.S. Centers for Disease Control and Prevention
Symptoms Testing U.S. Centers for Disease Control and Prevention
CDC Global Response to COVID-19_CDC Archive U.S. Centers for Disease Control and Prevention
CORD-19 COVID-19 Open Research Dataset (CORD-19), Allen Institute for AI
COVID-19-Global-Response-fact-sheet U.S. Centers for Disease Control and Prevention
FAQ COVID-19 Data and Surveillance_CDC Archive U.S. Centers for Disease Control and Prevention
Table 6: Knowledge bases categorized by source